# Long-time asymptotics of (1,3)-sign solitary waves for the damped nonlinear Klein-Gordon equation

### v1

@@ 1.6. Notation
Let $\{e_{1},\ldots,e_{d}\}$ denote the canonical basis of $\mathbb{R}^{d}$, and we denote ${\partial}_{k}$ as the derivative in the $x_{k}$-direction. The Euclidean inner product is denoted for any pair of vectors $x,y\in\mathbb{R}^{d}$ by
@@
In this paper, $X\lesssim Y$ means that $X\leq CY$ for some constant $C>0$. $X\sim Y$ means that $X\lesssim Y$ and $Y\lesssim X$. $X\ll Y$ means that there is a sufficiently small constant $c>0$ such that $X\leq cY$ The meaning of sufficiency should be clear from the context, or clarified later in more explicit form.
Moreover, in this paper we often deal with the second smallest value among three quantities. We therefore define
@@ 2. Preliminaries
In this section, we collect several preliminary results on the long-time dynamics of multi-solitary-wave solutions to (DNKG). Most of the material is taken from [21, Section2, Section3] and is included here for the reader’s convenience.
@@ 2.1. Basic properties of the ground state
Since $Q$ is the unique positive radial $H^{1}(\mathbb{R}^{d})$ solution to (1.3), we may write
@@
Furthermore, there exists a constant $c_{q}$ depending only on $d$ and $p$ such that
@@
Note that this definition is well defined. Furthermore, we have
@@
$\left|g(R+a)-g(R)\right|\lesssim\frac{g(R)}{R},$ (2.5)
@@ 2.2. Basic properties of the dynamics of $K$-solitary waves
We now summarize several results concerning the long-time dynamics of $K$-solitary waves. Let $\vec{u}$ be a $K$-solitary waves. Then there exist $\sigma\in\{-1,1\}^{K}$ and a function $z:[0,\infty)\to(\mathbb{R}^{d})^{K}$ such that (1.4) holds. For such $z$ and $\sigma$, we define
@@
In a multi-soliton decomposition, the choice of the center parameters $z_{k}(t)$ is not unique due to translation invariance. Following [21, Lemma 3.2], we fix this freedom by imposing orthogonality conditions that eliminate the contribution of the neutral directions associated with translations of the linearized operator around each soliton. This leads to a more convenient system for the modulation parameters.
@@
$\int\Bigl\{{\partial}_{t}u(t)+2\alpha\Bigl(u(t)-\sum_{k=1}^{K}\sigma_{k}Q(\cdot-z_{k}(t))\Bigr)\Bigr\}\,{\partial}_{l}Q(\cdot-z_{i}(t))=0.$ (2.9)
@@
Then, by [21, Section 3] $\mathcal{F}$ describes the leading-order interaction force between well-separated solitons.
@@
By Lemma 2.1, there exist $\sigma\in\{-1,1\}^{K}$ and a $C^{1}$ function $z:[0,\infty)\to(\mathbb{R}^{d})^{K}$ such that $\sigma$ and zzsatisfy (2.7), (2.8), and (2.9). Furthermore, by [21, Lemma 3.3], for $1<\theta<\min{(p-1,2)}$, we have
@@
In addition, by [21, Lemma 3.8] and
$e^{-\theta D}\ll\mathcal{F}(D),$
@@
Then, we introduce the result of $V$.
@@
$D(t)-\log{t}+\frac{d-1}{2}\log{(\log{t})}\lesssim 1$ (2.17)
$\mathcal{F}(D)\gtrsim t^{-1}$ (2.18)
@@
(2.16) represents [21, Lemma3.8]. (2.17) and (2.18) represent [21, Lemma3.9]. Thus, see[21, Lemma3.8, Lemma 3.9]. ∎
@@
then one can further improve the control on the remainder and the modulation parameters by [21, Lemma 3.11].
@@
3. Soliton interactions for $(1,3)$-sign solitary waves
Here, we consider $(1,3)$-sign solitary waves. First, to simplify the notation, we rewrite Lemma 2.1 in the following form.
@@ 3. Soliton interactions for $(1,3)$-sign solitary waves
$\int\{{\partial}_{t}u(t)+2\alpha(u(t)-Q(\cdot-z_{0}(t))+\sum_{k=1}^{3}Q(\cdot-z_{k}(t)))\}{\partial}_{l}Q(\cdot-z_{i}(t))=0.$ (3.3)
@@
In particular, by Lemma 3.1, wherever we assume in what follows that $\vec{u}$ is a $(1,3)$-sign solitary waves, we also take $z_{0},z_{1},z_{2},z_{3}$ to be $C^{1}$ functions satisfying (3.1), (3.2), and (3.3).
Here we introduce several definitions to further simplify the notation. First, in what follows we proceed using the functions $z_{k}$ above. Furthermore we define for $k=1,2,3$,
@@
is a Gram matrix. Hence, when we define $A$ as
$A(t)=1+2c_{12}c_{23}c_{13}-c_{12}^{2}-c_{13}^{2}-c_{23}^{2},$ (3.8)
we have for $t\geq 0$
@@
In this setting, we obtain the following estimates.
@@
Next, for $\rho_{1}$ we have
@@ 4. Estimates on interactions between like-signed solitons
From this section on, we analyze the behavior of the soliton centers. In the $(1,3)$-sign solitary waves setting, a total of six pairwise interactions are present, which makes it difficult to analyze the long-time dynamics of the solitons. We therefore begin by comparing the distances between like-signed solitons and those between oppositely signed solitons. To this end, we first introduce several definitions. First, we define $\tilde{D}$ and $\hat{D}$ as
@@ 4.1. Basic properties of $(1,3)$-sign solitary waves
$V=\mathcal{F}(\rho_{1})+\mathcal{F}(\rho_{2})+\mathcal{F}(\rho_{3})-\mathcal{F}(\rho_{12})-\mathcal{F}(\rho_{23})-\mathcal{F}(\rho_{13})\gg-\mathcal{F}(D).$ (4.4)
@@
Moreover, in this section we will sometimes need to compare the variation in the inter-soliton distances over time with the variation in the angles formed by the solitons. As an indicator for this purpose, we record the following.
@@
    $\left|c_{jk}(t)-c_{jk}(T)\right|\leq D^{-\epsilon}(T),$ (4.8)
    $\left|D(t)-D(T)\right|\lesssim D^{1-\epsilon}(T).$ (4.9)
@@
    Moreover,
    $\int_{T}^{T^{\prime}}\mathcal{F}(D(s))ds\sim D^{1-\epsilon}(T).$ (4.10)
@@
Fix $\epsilon>0$. First, by Lemma 3.3, there exists $C>1$ such that for $t\geq T$,
$|\dot{D}|\leq C\mathcal{F}(D),$ (4.11)
$|\dot{c}_{12}|+|\dot{c}_{13}|+|\dot{c}_{23}|\leq\frac{C\mathcal{F}(D)}{D}.$ (4.12)
@@
$\int_{T}^{T^{\prime}}\frac{\mathcal{F}(D(s))ds}{D(s)}=\frac{1}{C}D^{-\epsilon}(T).$ (4.13)
We show that, for this choice of $T^{\prime}$, (4.8), (4.9), and (4.10) hold.
First, we prove (4.8). To prove (4.8), fix $1\leq j<k\leq 3$ and take any $T\leq t\leq T^{\prime}$. Then by (4.12), we have
$|c_{jk}(t)-c_{jk}(T)|\leq\int_{T}^{t}|\dot{c}_{jk}(s)|ds\leq C\int_{T}^{T^{\prime}}\frac{\mathcal{F}(D(s))ds}{D(s)}=D^{-\epsilon}(T),$
which yields (4.8).
Next, we prove (4.9). For any $T\leq t\leq T^{\prime}$, we have
$\begin{aligned} \left|\log{D(t)}-\log{D(T)}\right|&=\left|\int_{T}^{t}\frac{\dot{D}(s)ds}{D(s)}\right|\\ &\leq\int_{T}^{t}\frac{|\dot{D}(s)|ds}{D(s)}\\ &\leq C\int_{T}^{T^{\prime}}\frac{\mathcal{F}(D(s))ds}{D(s)}=D^{-\epsilon}(T).\end{aligned}$ (4.14)
@@
$D(T)(1-2D^{-\epsilon}(T))\leq e^{-D^{-\epsilon}(T)}D(T),$
$e^{-D^{-\epsilon}(T)}D(T)\leq D(t)\leq e^{D^{-\epsilon}(T)}D(T),$
$e^{D^{-\epsilon}(T)}D(T)\leq D(T)(1+2D^{-\epsilon}(T)).$
Hence (4.9) follows.
Finally, we prove (4.10). By using (4.9), we have for $T\leq t\leq T^{\prime}$
$D(t)\sim D(T).$
Then by (4.13),
$D^{-\epsilon}(T)$ $\sim\int_{T}^{T^{\prime}}\frac{\mathcal{F}(D(s))ds}{D(s)}$
$\sim\frac{1}{D(T)}\int_{T}^{T^{\prime}}\mathcal{F}(D(s))ds,$
which yields (4.10). This completes the proof. ∎
@@
    $\left|\ |\varsigma_{1}-\varsigma_{2}|-M\ \right|$ $\lesssim M^{\frac{99}{100}},$
    $\left|\ |\varsigma_{1}-\varsigma_{3}|-M\ \right|$ $\lesssim M^{\frac{99}{100}},$
    $|\varsigma_{2}-\varsigma_{3}|-M$ $\gtrsim-M^{\frac{99}{100}},$
@@
    $\frac{\varsigma_{2}-\varsigma_{1}}{|\varsigma_{2}-\varsigma_{1}|}\cdot\frac{\varsigma_{3}-\varsigma_{1}}{|\varsigma_{3}-\varsigma_{1}|}-\frac{1}{2}$ $\lesssim M^{-\frac{1}{100}},$
    $\frac{\varsigma_{1}-\varsigma_{2}}{|\varsigma_{1}-\varsigma_{2}|}\cdot\frac{\varsigma_{3}-\varsigma_{2}}{|\varsigma_{3}-\varsigma_{2}|}-\frac{1}{2}$ $\gtrsim-M^{-\frac{1}{100}},$
    $\frac{\varsigma_{1}-\varsigma_{3}}{|\varsigma_{1}-\varsigma_{3}|}\cdot\frac{\varsigma_{2}-\varsigma_{3}}{|\varsigma_{2}-\varsigma_{3}|}-\frac{1}{2}$ $\gtrsim-M^{-\frac{1}{100}}.$
@@
    $\left|\ |\varsigma_{1}-\varsigma_{2}|-M\ \right|$ $\lesssim M^{\frac{99}{100}},$
    $|\varsigma_{1}-\varsigma_{3}|-M$ $\gtrsim-M^{\frac{99}{100}},$
    $|\varsigma_{2}-\varsigma_{3}|-M$ $\gtrsim-M^{\frac{99}{100}},$
@@
    $\frac{\varsigma_{1}-\varsigma_{3}}{|\varsigma_{1}-\varsigma_{3}|}\cdot\frac{\varsigma_{2}-\varsigma_{3}}{|\varsigma_{2}-\varsigma_{3}|}-\frac{1}{2}$ $\gtrsim-M^{-\frac{1}{100}}.$

### v2

@@ 1.6. Notation
Let $\{e_{1},\ldots,e_{d}\}$ denote the canonical basis of $\mathbb{R}^{d}$, and let ${\partial}_{k}$ denote the partial derivative with respect to $x_{k}$. The Euclidean inner product is denoted for any pair of vectors $x,y\in\mathbb{R}^{d}$ by
@@
In this paper, $X\lesssim Y$ means that $X\leq CY$ for some constant $C>0$. $X\sim Y$ means that $X\lesssim Y$ and $Y\lesssim X$. $X\ll Y$ means that there exists a sufficiently small constant $c>0$ such that $X\leq cY$. The meaning of sufficiency should be clear from the context, or clarified later in more explicit form. Moreover, in this paper we often deal with the second smallest value among three quantities. We therefore define
@@ 2. Preliminaries
In this section, we collect several preliminary results on the long-time dynamics of $K$-solitary waves solutions. Most of the material is taken from [20, Section 2 and 3] and is included here for the reader’s convenience.
@@ 2.1. Basic properties of the ground state
Since $Q$ is radial, there exists $q:[0,\infty)\to\mathbb{R}$ such that
@@
In addition, there exists a constant $c_{q}$ depending only on $d$ and $p$ such that
@@
We note that this definition is well-defined. Furthermore, we have
@@
$\left|g(R+a)-e^{-a}g(R)\right|\lesssim\frac{g(R)}{R},$ (2.5)
@@ 2.2. Basic properties of the dynamics of $K$-solitary waves
We now summarize several results concerning the long-time dynamics of $K$-solitary waves solutions. Let $\vec{u}$ be a $K$-solitary waves. Then there exist $\sigma\in\{-1,1\}^{K}$ and a function $z:[0,\infty)\to(\mathbb{R}^{d})^{K}$ such that (1.4) holds. For such $z$ and $\sigma$, we define
@@
In a multi-soliton decomposition, the choice of the center parameters $z_{k}(t)$ is not unique due to translation invariance. Following [20, Lemma 3.2], we fix this freedom by imposing orthogonality conditions that eliminate the contribution of the neutral directions associated with translations of the linearized operator around each soliton. This leads to a more convenient system for the modulation parameters.
@@
$\int_{\mathbb{R}^{d}}\Bigl\{{\partial}_{t}u(t)+2\alpha\Bigl(u(t)-\sum_{k=1}^{K}\sigma_{k}Q(\cdot-z_{k}(t))\Bigr)\Bigr\}\,{\partial}_{l}Q(\cdot-z_{i}(t))=0.$ (2.9)
@@
Then, by [20, Section 3] $\mathcal{F}$ describes the leading-order interaction force between well-separated solitons.
@@
By Lemma 2.1, there exist $\sigma\in\{-1,1\}^{K}$ and a $C^{1}$ function $z:[0,\infty)\to(\mathbb{R}^{d})^{K}$ such that $\sigma$ and $z$ satisfy (2.7), (2.8), and (2.9). Furthermore, by [20, Lemma 3.3], for $1<\theta<\min{(p-1,2)}$, we have
@@
In addition, by [20, Lemma 3.8] and
$e^{-\theta D}\lesssim\frac{\mathcal{F}(D)}{D},$
@@
We next record a useful estimate of $V$.
@@
$D(t)-\log{t}+\frac{d-1}{2}\log{(\log{t})}\lesssim 1,$ (2.17)
$\mathcal{F}(D)\gtrsim\frac{1}{t}.$ (2.18)
@@
(2.16) represents [20, Lemma 3.8]. (2.17) and (2.18) represent [20, Lemma 3.9]. Thus, see [20, Lemma 3.8, Lemma 3.9]. ∎
@@
then one can further improve the control on the remainder and the modulation parameters by [20, Lemma 3.11].
@@
3. Soliton interactions for $(1,3)$-sign solitary waves solutions
In this section, we consider $(1,3)$-sign solitary waves solutions. First, to simplify the notation, we rewrite Lemma 2.1 in the following form.
@@ 3. Soliton interactions for $(1,3)$-sign solitary waves solutions
$\int_{\mathbb{R}^{d}}\{{\partial}_{t}u(t)+2\alpha(u(t)-Q(\cdot-z_{0}(t))+\sum_{k=1}^{3}Q(\cdot-z_{k}(t)))\}{\partial}_{l}Q(\cdot-z_{i}(t))=0.$ (3.3)
@@
In particular, by Lemma 3.1, throughout the rest of this paper, whenever we assume that $\vec{u}$ is a $(1,3)$-sign solitary waves, we choose $C^{1}$ functions $z_{0},z_{1},z_{2},z_{3}$ so that (3.1), (3.2), and (3.3) hold.
Here we introduce several definitions to further simplify the notation. First, in what follows we proceed using the functions $z_{k}$ above. Furthermore, we define for $k=1,2,3$,
@@
is a Gram matrix. Hence, let $A(t)$ denote its determinant; that is,
$A(t)=1+2c_{12}c_{23}c_{13}-c_{12}^{2}-c_{13}^{2}-c_{23}^{2}.$ (3.8)
Then, we have for $t\geq 0$
@@
In this setting, we obtain the following identities.
@@
Next, we estimate $\rho_{1}$. By direct computation, we have
@@
By the above computations, if $\vec{u}$ is a $(1,3)$-sign solitary waves and satisfies (2.19), then Lemma 3.3 can be refined as in Lemma 2.4.
Corollary 3.4.
Let $\vec{u}$ be a $(1,3)$-sign solitary waves and suppose that it satisfies (2.19). Then for any $1<\theta<\min{(p-1,2)}$, the following hold.
  1. (1)
    $Z_{1},Z_{2},Z_{3}$ satisfy for $t\gg 1$
    $\begin{aligned} \dot{Z}_{1}&=2\mathcal{F}(\rho_{1})u_{1}+\mathcal{F}(\rho_{2})u_{2}+\mathcal{F}(\rho_{3})u_{3}\\ &\quad+\mathcal{F}(\rho_{12})u_{12}+\mathcal{F}(\rho_{13})u_{13}+O\left(e^{-\theta D}\right),\\ \dot{Z}_{2}&=\mathcal{F}(\rho_{1})u_{1}+2\mathcal{F}(\rho_{2})u_{2}+\mathcal{F}(\rho_{3})u_{3}\\ &\quad+\mathcal{F}(\rho_{21})u_{21}+\mathcal{F}(\rho_{23})u_{23}+O\left(e^{-\theta D}\right),\\ \dot{Z}_{3}&=\mathcal{F}(\rho_{1})u_{1}+\mathcal{F}(\rho_{2})u_{2}+2\mathcal{F}(\rho_{3})u_{3}\\ &\quad+\mathcal{F}(\rho_{31})u_{31}+\mathcal{F}(\rho_{32})u_{32}+O\left(e^{-\theta D}\right).\end{aligned}$ (3.21)
  2. (2)
    $\rho_{1},\ \rho_{2},\ \rho_{3}$ satisfy for $t\gg 1$
    $\begin{aligned} \dot{\rho_{1}}&=2\mathcal{F}(\rho_{1})+\mathcal{F}(\rho_{2})c_{12}+\mathcal{F}(\rho_{3})c_{13}\\ &\quad+\mathcal{F}(\rho_{12})u_{12}\cdot u_{1}+\mathcal{F}(\rho_{13})u_{13}\cdot u_{1}+O\left(e^{-\theta D}\right),\\ \dot{\rho}_{2}&=\mathcal{F}(\rho_{1})c_{12}+2\mathcal{F}(\rho_{2})+\mathcal{F}(\rho_{3})c_{23}\\ &\quad+\mathcal{F}(\rho_{21})u_{21}\cdot u_{2}+\mathcal{F}(\rho_{23})u_{23}\cdot u_{2}+O\left(e^{-\theta D}\right),\\ \dot{\rho}_{3}&=\mathcal{F}(\rho_{1})c_{13}+\mathcal{F}(\rho_{2})c_{23}+2\mathcal{F}(\rho_{3})\\ &\quad+\mathcal{F}(\rho_{31})u_{31}\cdot u_{3}+\mathcal{F}(\rho_{32})u_{32}\cdot u_{3}+O\left(e^{-\theta D}\right).\end{aligned}$ (3.22)
  3. (3)
    $c_{12},c_{13},c_{23}$ satisfy for $t\gg 1$
    $\begin{aligned} \dot{c}_{12}&=(1-c_{12}^{2})\left\{\frac{\mathcal{F}(\rho_{1})}{\rho_{2}}+\frac{\mathcal{F}(\rho_{2})}{\rho_{1}}+\frac{\mathcal{F}(\rho_{12})}{\rho_{12}}\left(\frac{\rho_{2}}{\rho_{1}}+\frac{\rho_{1}}{\rho_{2}}\right)\right\}\\ &\quad+\mathcal{F}(\rho_{3})\left(\frac{c_{23}-c_{13}c_{12}}{\rho_{1}}+\frac{c_{13}-c_{23}c_{12}}{\rho_{2}}\right)\\ &\quad+\frac{\mathcal{F}(\rho_{13})}{\rho_{13}}\frac{\rho_{3}(c_{23}-c_{12}c_{13})}{\rho_{1}}+\frac{\mathcal{F}(\rho_{23})}{\rho_{23}}\frac{\rho_{3}(c_{13}-c_{12}c_{23})}{\rho_{2}}+O\left(e^{-\theta D}\right),\\ \dot{c}_{13}&=(1-c_{13}^{2})\left\{\frac{\mathcal{F}(\rho_{1})}{\rho_{3}}+\frac{\mathcal{F}(\rho_{3})}{\rho_{1}}+\frac{\mathcal{F}(\rho_{13})}{\rho_{13}}\left(\frac{\rho_{3}}{\rho_{1}}+\frac{\rho_{1}}{\rho_{3}}\right)\right\}\\ &\quad+\mathcal{F}(\rho_{2})\left(\frac{c_{23}-c_{12}c_{13}}{\rho_{1}}+\frac{c_{12}-c_{13}c_{23}}{\rho_{3}}\right)\\ &\quad+\frac{\mathcal{F}(\rho_{12})}{\rho_{12}}\frac{\rho_{2}(c_{23}-c_{12}c_{13})}{\rho_{1}}+\frac{\mathcal{F}(\rho_{23})}{\rho_{23}}\frac{\rho_{2}(c_{12}-c_{13}c_{23})}{\rho_{3}}+O\left(e^{-\theta D}\right),\\ \dot{c}_{23}&=(1-c_{23}^{2})\left\{\frac{\mathcal{F}(\rho_{2})}{\rho_{3}}+\frac{\mathcal{F}(\rho_{3})}{\rho_{2}}+\frac{\mathcal{F}(\rho_{23})}{\rho_{23}}\left(\frac{\rho_{3}}{\rho_{2}}+\frac{\rho_{2}}{\rho_{3}}\right)\right\}\\ &\quad+\mathcal{F}(\rho_{1})\left(\frac{c_{13}-c_{12}c_{23}}{\rho_{2}}+\frac{c_{12}-c_{13}c_{23}}{\rho_{3}}\right)\\ &\quad+\frac{\mathcal{F}(\rho_{12})}{\rho_{12}}\frac{\rho_{1}(c_{13}-c_{12}c_{23})}{\rho_{2}}+\frac{\mathcal{F}(\rho_{13})}{\rho_{13}}\frac{\rho_{1}(c_{12}-c_{13}c_{23})}{\rho_{3}}+O\left(e^{-\theta D}\right).\end{aligned}$ (3.23)
  4. (4)
    $\rho_{12},\ \rho_{13},\ \rho_{23}$ satisfy for $t\gg 1$
    $\begin{aligned} \dot{\rho}_{12}&=-2\mathcal{F}(\rho_{12})+\mathcal{F}(\rho_{2})(u_{12}\cdot u_{2})+\mathcal{F}(\rho_{1})(u_{1}\cdot u_{21})\\ &\quad-\mathcal{F}(\rho_{23})(u_{21}\cdot u_{23})-\mathcal{F}(\rho_{13})(u_{12}\cdot u_{13})+O\left(e^{-\theta D}\right),\\ \dot{\rho}_{13}&=-2\mathcal{F}(\rho_{13})+\mathcal{F}(\rho_{1})(u_{31}\cdot u_{1})+\mathcal{F}(\rho_{3})(u_{13}\cdot u_{3})\\ &\quad-\mathcal{F}(\rho_{12})(u_{13}\cdot u_{12})-\mathcal{F}(\rho_{23})(u_{31}\cdot u_{32})+O\left(e^{-\theta D}\right),\\ \dot{\rho}_{23}&=-2\mathcal{F}(\rho_{23})+\mathcal{F}(\rho_{2})(u_{32}\cdot u_{2})+\mathcal{F}(\rho_{3})(u_{23}\cdot u_{3})\\ &\quad-\mathcal{F}(\rho_{12})(u_{21}\cdot u_{23})-\mathcal{F}(\rho_{13})(u_{32}\cdot u_{31})+O\left(e^{-\theta D}\right).\end{aligned}$ (3.24)
@@ 4. Estimates on interactions between like-signed solitons
In this section and what follows, we analyze the behavior of the soliton centers. In the $(1,3)$-sign configuration, there are six pairwise interactions, which makes the analysis of the long-time soliton dynamics more delicate. We therefore begin by comparing the distances between solitons of the same sign and those between solitons of opposite signs. To this end, we first introduce several definitions. First, we define $\tilde{D}$ and $\hat{D}$ as
@@ 4.1. Basic properties of $(1,3)$-sign solitary waves
$V=\mathcal{F}(\rho_{1})+\mathcal{F}(\rho_{2})+\mathcal{F}(\rho_{3})-\mathcal{F}(\rho_{12})-\mathcal{F}(\rho_{23})-\mathcal{F}(\rho_{13})>-\frac{1}{100}\mathcal{F}(D).$ (4.4)
@@
In this section, we need to analyze delicate variations in the inter-soliton distances. To this end, we establish the following lemma concerning $D$ and $\mathcal{F}(D)$.
@@
    $\left|D(t)-D(T)\right|\leq 2D^{\epsilon}(T).$ (4.7)
@@
    Moreover, there exists $0<c_{\bullet}<1$ depending on $\vec{u}$ such that
    $c_{\bullet}D^{\epsilon}(T)\leq\int_{T}^{T^{\prime}}\mathcal{F}(D(s))ds\leq\frac{1}{c_{\bullet}}D^{\epsilon}(T).$ (4.8)
@@
Fix $\epsilon>0$. First, by Lemma 3.3, there exists $C>1$ depending on $\vec{u}$ such that for $t\geq T$
$|\dot{D}|\leq C\mathcal{F}(D).$ (4.9)
@@
$\int_{T}^{T^{\prime}}\frac{\mathcal{F}(D(s))ds}{D(s)}=\frac{1}{C}D^{\epsilon-1}(T).$ (4.10)
We show that, for this choice of $T^{\prime}$, (4.7), and (4.8) hold.
First, we prove (4.7). For any $T\leq t\leq T^{\prime}$, we have
$\begin{aligned} \left|\log{D(t)}-\log{D(T)}\right|&=\left|\int_{T}^{t}\frac{\dot{D}(s)ds}{D(s)}\right|\\ &\leq\int_{T}^{t}\frac{|\dot{D}(s)|ds}{D(s)}\\ &\leq C\int_{T}^{T^{\prime}}\frac{\mathcal{F}(D(s))ds}{D(s)}=D^{\epsilon-1}(T).\end{aligned}$ (4.11)
@@
$D(T)(1-2D^{\epsilon-1}(T))\leq e^{-D^{\epsilon-1}(T)}D(T),$
$e^{-D^{\epsilon-1}(T)}D(T)\leq D(t)\leq e^{D^{\epsilon-1}(T)}D(T),$
$e^{D^{\epsilon-1}(T)}D(T)\leq D(T)(1+2D^{\epsilon-1}(T)),$
which implies (4.7).
Next, we prove (4.8). By (4.7), we have for $T\leq t\leq T^{\prime}$
$\frac{1}{2}D(T)\leq D(t)\leq 2D(T).$
Then by (4.10), we have
$D^{\epsilon-1}(T)$ $=C\int_{T}^{T^{\prime}}\frac{\mathcal{F}(D(s))ds}{D(s)}$
$\leq\frac{2C}{D(T)}\int_{T}^{T^{\prime}}\mathcal{F}(D(s))ds,$
and
$D^{\epsilon-1}(T)$ $=C\int_{T}^{T^{\prime}}\frac{\mathcal{F}(D(s))ds}{D(s)}$
$\geq\frac{C}{2D(T)}\int_{T}^{T^{\prime}}\mathcal{F}(D(s))ds,$
which yields (4.8) by choosing $c_{\bullet}=\min{(\frac{1}{2C},\frac{C}{2})}$. This completes the proof. ∎
@@
    $\left|\ |\varsigma_{1}-\varsigma_{2}|-M\ \right|$ $\leq M^{\frac{99}{100}},$
    $\left|\ |\varsigma_{1}-\varsigma_{3}|-M\ \right|$ $\leq M^{\frac{99}{100}},$
    $|\varsigma_{2}-\varsigma_{3}|-M$ $\geq-M^{\frac{99}{100}},$
@@
    $\frac{\varsigma_{2}-\varsigma_{1}}{|\varsigma_{2}-\varsigma_{1}|}\cdot\frac{\varsigma_{3}-\varsigma_{1}}{|\varsigma_{3}-\varsigma_{1}|}-\frac{1}{2}$ $\leq 5M^{-\frac{1}{100}},$
    $\frac{\varsigma_{1}-\varsigma_{2}}{|\varsigma_{1}-\varsigma_{2}|}\cdot\frac{\varsigma_{3}-\varsigma_{2}}{|\varsigma_{3}-\varsigma_{2}|}-\frac{1}{2}$ $\geq-5M^{-\frac{1}{100}},$
    $\frac{\varsigma_{1}-\varsigma_{3}}{|\varsigma_{1}-\varsigma_{3}|}\cdot\frac{\varsigma_{2}-\varsigma_{3}}{|\varsigma_{2}-\varsigma_{3}|}-\frac{1}{2}$ $\geq-5M^{-\frac{1}{100}}.$
@@
    $\left|\ |\varsigma_{1}-\varsigma_{2}|-M\ \right|$ $\leq M^{\frac{99}{100}},$
    $|\varsigma_{1}-\varsigma_{3}|-M$ $\geq-M^{\frac{99}{100}},$
    $|\varsigma_{2}-\varsigma_{3}|-M$ $\geq-M^{\frac{99}{100}},$
@@
    $\frac{\varsigma_{1}-\varsigma_{3}}{|\varsigma_{1}-\varsigma_{3}|}\cdot\frac{\varsigma_{2}-\varsigma_{3}}{|\varsigma_{2}-\varsigma_{3}|}-\frac{1}{2}$ $\geq-5M^{-\frac{1}{100}}.$
Remark 4.1.
The lemma depends only on the configuration of three points and is therefore essentially two-dimensional. Thus, to gain geometric intuition, one may restrict attention to a planar triangle, in which the claim is most easily visualized.